OpenStudy (anonymous):

sqrt{45} sqrt{60} sqrt{45} sqrt{60} @Mathematics

OpenStudy (anonymous):

\[\sqrt{45} \times \sqrt{60}\]

OpenStudy (anonymous):

9/12

OpenStudy (lilg132):

51.961524227066318805823390245176

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

\[3\sqrt{5}+2\sqrt{15}\]=51.9615

OpenStudy (anonymous):

\[\sqrt{9*5} \times \sqrt{4*5*3} = 30\sqrt{3}\]

OpenStudy (anonymous):

ok so many answers lol